OpenStudy (anonymous):

Calculte the molarity: 40.5g of calcium chloride in 600.0 mL of solution

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) 40.5/(40.1)+(35.5*2)=0.364 Number of Moles=Concentration(mol/dm^-3)*volume(dm^3) To Change ml to dm^3 we divide by 1000 0.364=600/1000*Concentration 600(Concentration)=1000*0.364 Concentration=1000*0.364/600=0.6075 mol/dm^3

OpenStudy (anonymous):

Thank you!!!